# Asymptotic confidence bands for the histogram regression estimator

### v1

@@ Abstract
In a multivariate nonparametric regression model with a Hölder continuous regression function and heteroscedastic noise asymptotic uniform confidence bands are constructed based on the histogram estimator. The radius of the confidence bands does not depend on an extreme value distribution, but instead can be numerically calculated for the chosen binning.
Keywords: Multivariate nonparametric regression, histogram, confidence bands, heteroscedastic noise
@@ 1 Introduction
We consider classical histogram estimators for the regression function $m$ in a nonparametric regression model $Y=m(X)+\varepsilon$ with multivariate covariates $X$ and heteroscedastic noise, and derive asymptotic confidence bands.
@@
There is a substantial amount of literature on confidence bands based on different nonparametric regression estimators. Johnston, (1982) has constructed confidence bands for the Nadaraya-Watson kernel estimator. Härdle, (1989) has shown asymptotic uniform confidence bands for a wider class of regression estimators, the M-smoothers. Typically, confidence bands rely on an undersmoothing of the estimator to reduce the bias relative to the stochastic error. An alternative is a direct bias correction, which is used, for instance, by Eubank and Speckman, (1993), who have considered a deterministic, uniform design for local constant regression estimation, and by Xia, (1998) considering a random design under dependence and local linear estimation. There are also bootstrap confidence bands for nonparametric regression, see for instance Hall, (1993), Neumann and Polzehl, (1998), Claeskens and Van Keilegom, (2003) or Hall and Horowitz, (2013). Sabbah, (2014) has shown confidence bands for quantile estimators which is an alternative to a regression model, and Birke et al., (2010) have proved confidence bands in an inverse regression model.
@@
The aforementioned articles only considered univariate regression. In the multivariate case Konakov and Piterbarg, (1984) have analyzed the asymptotic distribution of the maximal deviation for the Nadaraya-Watson estimate in a random design setting. Proksch, (2016) has proved confidence bands in a multivariate regression model with fixed, deterministic design for a general class of estimators that includes, for example, local polynomial estimators. Chao et al., (2017) have covered confidence bands for multivariate quantile and expectile regression with mean regression as a special case.
Despite the popularity and widespread application of histogram estimators, the only uniform confidence band result for a histogram estimator that we are aware of is by Smirnov, (1950) and covers univariate density estimation. We want to close this gap by presenting a confidence band result in a multivariate regression model.
@@
In an undersmooting regime we construct an honest asymptotic uniform confidence band in the class of $\alpha$-Hölder regular regression functions for $\alpha\in(0,1]$. The diameter of our confidence band is naturally determined by the locally constant variance of the histogram estimator together with quantiles of the distribution of the maximum $\max_{j=1,\dots,J}|Z_{j}|$ of independent standard normal random variables $(Z_{j})_{j\geq 1}$ where $J$ is given by the number of cells in the partition of the histogram. In particular, our confidence bands do not rely on an extreme value distribution. We also incorporate the estimation of the unknown variance and a possible unknown distribution of the covariates.
@@
We obtain a confidence band result for the simple case of a Hölder continuous regression function for every covariate dimension, without a continuity assumption on the covariate density. In contrast for smooth nonparametric regression estimators typically it is assumed that for increasing covariate dimension the number of existing derivatives of the regression function and the covariate density have to increase, see Chao et al., (2017), for instance. Instead we apply error moment existence, but the assumptions are not restrictive in the sense of exponential moments.
@@
A central part of the proof for the histogram based confidence band is the application of a result by Chernozhukov et al., 2014b regarding the Gaussian approximation for the suprema of empirical processes. Using this direct approximation of the supremum, it is not necessary to approximate a whole empirical process uniformly. This method has also been applied by Patschkowski and Rohde, (2019) to construct adaptive confidence bands for probability densities. It is an important difference to previous results in the literature like those by Johnston, (1982), Claeskens and Van Keilegom, (2003) or Chao et al., (2017), and the proof method might be extended to different estimators.
@@
The main result is presented in Section 2. The proof strategy is outlined in Section 3, while detailed proofs are postponed to Section 4.
@@ 2 Histogram estimator and confidence bands
On some probability space $(\Omega,\mathcal{A},\mathbb{P})$ we observe the training sample $(X_{i},Y_{i})_{i=1}^{n}$ consisting of $n$ observations which are independent copies of the random variable $(X,Y)$ with $X\in[0,1]^{p}$ and $Y\in\mathbb{R}$. The goal is to estimate the regression function $m\colon[0,1]^{p}\to\mathbb{R}$ with
@@
The histogram estimator relies on a local constant approximation of $m$. For a given partition of the feature space $0,1]^{p}$ the regression function is estimated by the cell-wise average of the response variables in each cell of the partition. We will choose an equidistant grid to construct the partition such that all cells are hypercubes of the same size. For $\delta^{-1}\in\mathbb{N}$ we consider the grid with edge length $\delta$ on $0,1]^{p}$. Here $\delta=\delta_{n}$ is a sequence depending on the sample size with $\delta\to 0$ for $n\to\infty$. For $x\in[0,1]^{p}$ let $A_{\delta}(x)$ be the hypercube with volume $\delta^{p}$ that contains $x$. Note that the equidistant grid is considered for simplicity and all results carry over to more irregular partitions as long as the volume and the diameters of all cells in the same order. The histogram estimator is then defined by
$\hat{m}(x)=\sum_{j=1}^{n}Y_{j}\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}$
where $\mathbb{I}\{\cdot\}$ denotes the indicator function. If there is no observation in $A_{\delta}(x)$ we define the estimator as zero on this cell.
In the following theorem we present a confidence band result based on the histogram estimator. Let $\mathcal{H}^{\alpha}(C_{H})$ denote the set of all bounded $\alpha$-Hölder functions $0,1]^{p}\to\mathbb{R}$ with Hölder regularity $\alpha\in(0,1]$ and both Hölder constant and supremum norm bounded by $C_{H}>0$.
@@
Let $\alpha\in(0,1],C_{H}>0$ and assume that
$n\delta^{2\alpha+p}(\log n)^{2}\to 0.$ (3)
Further let $4\leq\nu\in\mathbb{N}$ be such that $\mathbb{E}\left[|\varepsilon|^{\nu}\right]<\infty$ and
$\frac{(\log n)^{5}}{\delta^{p}n^{1-2/\nu}}$ $\to 0$ (4)
and assume that $\sup_{x\in[0,1]^{p}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$. For independent standard normally distributed random variables $(Z_{j})_{j\in\mathbb{N}}$ and some $\beta\in(0,1)$ let $c_{\delta}(\beta)>0$ be given by
$\mathbb{P}\left(\max_{j=1,\ldots,\delta^{-p}}|Z_{j}|\leq c_{\delta}(\beta)\right)=1-\beta.$
@@
$\mathcal{C}_{n}(x)=\left[\hat{m}(x)-c_{\delta}(\beta)(\tau_{\delta}(x)n)^{-1/2},\hat{m}(x)+c_{\delta}(\beta)(\tau_{\delta}(x)n)^{-1/2}\right]$
@@
$\liminf_{n\to\infty}\inf_{m\in\mathcal{H}^{\alpha}(C_{H})}\mathbb{P}\left(m(x)\in\mathcal{C}_{n}(x),\,\forall x\in[0,1]^{p}\right)\geq 1-\beta.$
@@
$\max_{j=1,\ldots,\delta^{-p}}|Z_{j}|=\mathcal{O}_{\mathbb{P}}(\sqrt{2p\log\delta^{-1}})=\mathcal{O}_{\mathbb{P}}(\sqrt{\log n})$ (6)
due to $\log\delta^{-1}\lesssim\log n$ which is implied by assumption (4). Therefore, $c_{\delta}(\beta)\lesssim\sqrt{\log n}$.
@@
By (1) and (2) we have
$c_{X}\delta^{p}\leq p_{\delta}(x)\leq C_{X}\delta^{p}$ (7)
@@
$\frac{c_{X}}{C_{\sigma^{2}}}\delta^{p}\leq\tau_{\delta}(x)\leq\frac{C_{X}}{c_{\sigma^{2}}}\delta^{p}.$ (8)
Therefore, the diameter of the confidence band $\mathcal{C}_{n}$ is of order $\sqrt{(\log n)/(n\delta^{p})}$ uniformly in $x\in[0,1]^{p}$. In the homoscedastic case with constant $\sigma^{2}(\cdot)\equiv\sigma^{2}$ we obtain $\tau_{\delta}(x)=p_{\delta}(x)/\sigma^{2}$. For uniformly distributed covariates on $0,1]^{p}$ we have $p_{\delta}(x)=\delta^{p}$ and in that case the confidence band simplifies to
$\mathcal{C}_{n}(x)=\Big{[}\hat{m}(x)-c_{\delta}(\beta)\frac{\sigma}{\sqrt{n\delta^{p}}},\hat{m}(x)+c_{\delta}(\beta)\frac{\sigma}{\sqrt{n\delta^{p}}}\Big{]}.$
@@
$n\delta^{p+2\alpha}\to 0\text{ and }n\delta^{p}/(n^{2/\nu})\to\infty,$
where $2/\nu$ is arbitrary small if all moments of $\varepsilon$ exist. Since the approximation error of the histogram estimator is of order $\delta^{\alpha}$ while the stochastic error term is of order $(n\delta^{p})^{-1/2}$, the first assumption is an undersmoothing condition. The second condition implies that the average number of observations in all cells goes to infinity such that the estimator is consistent. As $p$ increases, the assumptions become more restrictive which is the common curse of dimensionality problem for nonparametric function estimation. However, the assumptions are not contradictory, and for every dimension $p$ no bias reduction is needed. In contrast, the confidence bands by Chao et al., (2017) require increasing smoothness orders for the regression function $m$ and covariate density $f_{X}$ if the dimension $p$ increases.
We are not aware of any other confidence band result in the literature for the histogram estimator in the multivariate or univariate regression case. However, the structure of the histogram with fixed deterministic cells might imply that the difference in the proof from the univariate to the multivariate case is minor. The first result for the histogram by Smirnov, (1950) studies univariate density estimation. In addition to the different estimation problem, this result differs from our result in another crucial aspect. The asymptotic bands rely on an extreme value distribution that is independent of $\delta$. Instead our result provides bands with explicit dependence on $\delta$. Thus, the knowledge of the limit distribution is not necessary, and the accuracy of the bands does not depend on the rate of convergence to that limit distribution. Applying an approximation by the maximum of Gaussian random variables instead of a limit distribution is also suggested by Kreiß et al., (2025) for confidence bands for the spectral density in time series analysis.
From Theorem 2.1 we obtain a uniform rate of convergence of the histogram estimator.
@@
Assume $\sup_{x\in[0,1]^{p}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$ and $\mathbb{E}\left[|\varepsilon|^{\nu}\right]<\infty$ for some $4\leq\nu\in\mathbb{N}$ such that $(\log n)^{5}/(\delta^{p}n^{1-2/\nu})\to 0$. Then we have
$\sup_{x\in[0,1]^{p}}|\hat{m}(x)-m(x)|=\mathcal{O}_{\mathbb{P}}\left(\delta^{\alpha}+\sqrt{\frac{\log n}{n\delta^{p}}}\right).$
For $\nu>2+p/\alpha$ and $\delta=(n/\log n)^{-1/(2\alpha+p)}$ we obtain the uniform rate of convergence $\mathcal{O}_{\mathbb{P}}((n/\log n)^{-\alpha/(2\alpha+p)})$.
@@
We will now consider the case where the covariate distribution and the variance function of the noise are unknown. To this end, let $\hat{\tau}_{\delta}(x)$ be an estimator for $\tau_{\delta}(x)$ from (5) and define the confidence band as
$\hat{\mathcal{C}}_{n}(x)=\left[\hat{m}(x)-c_{\delta}(\beta)(\hat{\tau}_{\delta}(x)n)^{-1/2},\hat{m}(x)+c_{\delta}(\beta)(\hat{\tau}_{\delta}(x)n)^{-1/2}\right].$
@@
Under the assumptions of Theorem 2.1 and
$\sup_{x\in[0,1]^{p}}\left|\frac{\hat{\tau}_{\delta}(x)}{\tau_{\delta}(x)}-1\right|=o_{\mathbb{P}}((\log n)^{-3/2})$
@@
$\liminf_{n\to\infty}\inf_{m\in\mathcal{H}^{\alpha}(C_{H})}\mathbb{P}\left(m(x)\in\hat{\mathcal{C}}_{n}(x),\,\forall x\in[0,1]^{p}\right)\geq 1-\beta.$
The proof is again postponed to Section 4. To construct the estimator $\hat{\tau}_{\delta}(x)$, we first consider the homoscedastic case where $\tau_{\delta}(x)=p_{\delta}(x)/\sigma^{2}$ can be estimated by $\hat{\tau}_{\delta}(x)=\hat{p}_{\delta}(x)/\hat{\sigma}^{2}$ with
$\hat{p}_{\delta}(x)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}.$ (9)
Note that $\hat{p}_{\delta}(x)$ is the density histogram estimator for the density function $f_{X}$ of the covariate distribution. This estimator has the desired rate which we show in Lemma 4.2. The constant variance can be estimated using residuals based on the histogram regression estimator, i.e.
$\hat{\sigma}^{2}=\frac{1}{n}\sum_{i=1}^{n}\big{(}Y_{i}-\hat{m}(X_{i})\big{)}^{2}.$

### v2

# arXiv:2508.12391v2[math.ST] 27 Feb 2026

@@ Abstract
Asymptotic uniform confidence bands are constructed for a multivariate nonparametric regression model with heteroscedastic noise, employing histogram estimators under flexible partition conditions. The construction is especially applicable to unsmooth regression functions of Hölder regularity less than one. While the radius of the confidence bands could be approximated via the Gumbel distribution, our construction does not depend on an extreme value distribution, but instead can be explicitly calculated for the chosen partition.
Keywords: Multivariate nonparametric regression, partitioning estimator, confidence bands, heteroscedastic noise
@@ 1 Introduction
We consider histogram estimators for the regression function $m$ in a nonparametric regression model $Y=m(X)+\varepsilon$ with multivariate covariates $X$ and heteroscedastic noise. Our aim is to derive asymptotic histogram based confidence bands.
@@
There is a substantial amount of literature on confidence bands based on different nonparametric regression estimators. Johnston, (1982) has constructed confidence bands for the Nadaraya-Watson kernel estimator. Härdle, (1989) has shown asymptotic uniform confidence bands for a wider class of regression estimators, the M-smoothers. Typically, confidence bands rely on an undersmoothing of the estimator to reduce the bias relative to the stochastic error. An alternative is a direct bias correction, which is used, for instance, by Eubank and Speckman, (1993), who have considered a deterministic, uniform design for local constant regression estimation, and by Xia, (1998) considering a random design under dependence and local linear estimation. There are also bootstrap confidence bands for nonparametric regression, see for instance Hall, (1993), Neumann and Polzehl, (1998), Claeskens and Van Keilegom, (2003) or Hall and Horowitz, (2013). Based on spline estimators regression confidence bands were constructed by Wang and Yang, (2009) in the random covariate case, and by Krivobokova et al., (2010) in the deterministic design case. Cai et al., (2019) considered confidence bands for the mean and the variance function for fixed design. These confidence bands are based on Nadaraya-Watson-type estimators, where for the variance case the residuals are estimated based on spline-estimators. Sabbah, (2014) has shown confidence bands for quantile estimators which is an alternative to a regression model, and Birke et al., (2010) have proved confidence bands in an inverse regression model.
@@
The aforementioned articles only considered univariate regression. For multidimensional covariates often semiparametric regression models are considered due to the curse of dimensionality. For instance, confidence bands for the link function in a single-index model were developed by Gu and Yang, (2015), for additive regression functions by Härdle et al., (2015), and for additive semi-parametric regression functions by Cattaneo et al., (2024). In contrast we focus on multivariate nonparametric regression models without additional structural assumptions. In the multivariate case Konakov and Piterbarg, (1984) have analyzed the asymptotic distribution of the maximal deviation for the Nadaraya-Watson estimate in a random design setting. Proksch, (2016) has proved confidence bands in a multivariate regression model with fixed, deterministic design for a general class of estimators that includes, for example, local polynomial estimators. Chao et al., (2017) have covered confidence bands for multivariate quantile and expectile regression with mean regression as a special case.
Due to their easy and efficient implementation and clear interpretability histogram regression estimators (also called regressogram) enjoy a high popularity and widespread application, see e.g. Györfi et al., (2002); Wasserman, (2006). Also in the development of advanced statistical methodologies, histogram estimators continue to serve as a fundamental building block, e.g. for regression trees Breiman et al., (1984) or in the context of privacy, see Berrett et al., (2021). In a density estimation setting the study of confidence bands goes back to Smirnov, (1950). The regression histogram estimators are the simplest case of partitioning-based estimators whose asymptotic properties are studied in depth by Cattaneo et al., (2020). The latter article also includes confidence bands with a focus on general (bias corrected) partitioning-based estimators. For the case of multivariate histogram estimators the applicability of their construction is however limited. We want to close this gap by presenting an easy to implement confidence band result in a multivariate regression model for histogram estimators under very mild conditions on the partition. This especially allows us to incorporate a-priori knowledge on the regression function into the construction of the partition to improve the finite sample performance.
@@
In an undersmoothing regime we construct an honest asymptotic uniform confidence band in the class of $\alpha$-Hölder regular regression functions for $\alpha\in(0,1]$. Note that all above mentioned articles on confidence bands require the regression function to be at least continuously differentiable while we will allow for regression functions which are less than Lipschitz regular. The diameter of our confidence band is naturally determined by the locally constant variance of the histogram estimator together with quantiles of the distribution of the maximum $\max_{j=1,\dots,\Delta}|Z_{j}|$ of independent standard normal random variables $(Z_{j})_{j\geq 1}$ where $\Delta$ is given by the number of cells of the partition of the histogram. Whereas our confidence intervals do not need to be based on an extreme value distribution, our critical values can be related to the quantiles of the Gumbel distribution. We also incorporate the estimation of the unknown variance and a possible unknown distribution of the covariates.
@@
We obtain a confidence band result for the simple case of a Hölder continuous regression function for every covariate dimension, without a continuity assumption on the covariate density. In contrast, for smooth nonparametric regression estimators it is typically assumed that for increasing covariate dimension the number of existing derivatives of the regression function and the covariate density has to increase, see, for instance, Chao et al., (2017). Instead we only exploit finite higher moments of the error distribution, an assumption which is especially satisfied for finite exponential moments.
@@
A central part of the proof for the histogram based confidence band is the application of a result by Chernozhukov et al., 2014b regarding the Gaussian approximation for the suprema of empirical processes. Using this direct approximation of the supremum, it is not necessary to approximate a whole empirical process uniformly. This method has also been applied by Patschkowski and Rohde, (2019) to construct adaptive confidence bands for probability densities. It is a key difference to previous results in the literature like those by Johnston, (1982), Claeskens and Van Keilegom, (2003) or Chao et al., (2017), and the proof method might be extended to different estimators.
@@
The main result is presented and discussed in Section 2. The proof strategy is outlined in Section 3. In Section 4 we show a numerical illustration. Some conclusions are given in Section 5. Detailed proofs are postponed to Section 6 and in Appendix A extensions are discussed.
@@ 2 Histogram estimator and confidence bands
On some probability space $(\Omega,\mathcal{A},\mathbb{P})$ we observe the training sample $(X_{i},Y_{i})_{i=1}^{n}$ consisting of $n$ observations which are independent copies of the random variable $(X,Y)$ with $X\in\mathcal{X}$, $Y\in\mathbb{R}$ and where $\mathcal{X}\subseteq\mathbb{R}^{p}$ is some bounded feature space. The goal is to estimate the regression function $m\colon\mathcal{X}\to\mathbb{R}$ with
@@
The histogram estimator relies on a local constant approximation of $m$. For a given partition of the feature space $\mathcal{X}$ the regression function is estimated by the cell-wise average of the response variables in each cell of the partition. We will impose the following regularity assumption on the partition:
Assumption 2.1.
Let $\bigcup_{k=1}^{\Delta}A_{k}=\mathcal{X}$ be a partition of $\mathcal{X}$ of size $\Delta=\Delta_{n}\in\mathbb{N}$ into disjoint subsets $A_{k}\subset\mathcal{X},k=1,\dots,\Delta$, satisfying for some constants $c_{p},C_{p}>0$ that
$c_{p}\Delta^{-1}\leq\operatorname{vol}(A_{k})\leq C_{p}\Delta^{-1}\text{ and }\operatorname{diam}(A_{k})\leq C_{p}\Delta^{-1/p}\text{ for all }k=1,\dots,\Delta$ (3)
with volume $\operatorname{vol}(A_{k})$ and diameter $\operatorname{diam}(A_{k}):=\sup_{x_{1},x_{2}\in A_{k}}\|x_{1}-x_{2}\|$ of $A_{k}$.
On the one hand, the assumption guarantees that the overall volume $\operatorname{vol}(\mathcal{X})$ is roughly equally distributed over the cells of the partition. The condition on diameter ensures that the cells of the partition do not degenerate in any direction, but instead spread similarly in all directions. On the other hand, Assumption 2.1 allows for a flexible choice of the partition such that, e.g. prior knowledge of the regression function could be incorporated in the construction to improve the finite sample performance. In Section 4 we will see a numerical example in this direction.
For $x\in\mathcal{X}$ let $A_{\Delta}(x)$ be the subset $A_{k}$ that contains $x$. The histogram estimator is then defined by
$\widehat{m}(x)=\sum_{j=1}^{n}Y_{j}\frac{\mathbb{I}\{X_{j}\in A_{\Delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}$ (4)
where $\mathbb{I}\{\cdot\}$ denotes the indicator function. If there is no observation in $A_{\Delta}(x)$ we define the estimator as zero on this cell. The simplicity of histogram estimators comes with some computational advantages. First, they have lower computational costs than kernel estimators when many evaluations are needed ($O((m\wedge\Delta)n)$ instead of $O(mn)$ for $m$ point evaluations). Second, they serve as summary statistics to compress data and reduce storage space from $n$ observations to only $\Delta$ cell-wise means.
In the following theorem we present a confidence band result based on the histogram estimator. Let $\mathcal{H}^{\alpha}(C_{H})$ denote the set of all bounded $\alpha$-Hölder functions $m\colon\mathcal{X}\to\mathbb{R}$ with Hölder regularity $\alpha\in(0,1]$ and both Hölder constant and supremum norm bounded by $C_{H}>0$.
@@
Let the partition satisfy Assumption 2.1. Let $\alpha\in(0,1],C_{H}>0$ and assume that
$n\Delta^{-1-2\alpha/p}(\log n)^{2}\to 0.$ (5)
Assume that $\sup_{x\in\mathcal{X}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$ and $\mathbb{E}\left[|\varepsilon|^{\nu}\right]<\infty$ for some $4\leq\nu\in\mathbb{N}$ such that
$\frac{\Delta(\log n)^{5}}{n^{1-2/\nu}}$ $\to 0.$ (6)
For independent standard normally distributed random variables $(Z_{j})_{j\in\mathbb{N}}$ and some $\beta\in(0,1)$ let $c_{\Delta}(\beta)>0$ be given by
$\mathbb{P}\left(\max_{j=1,\ldots,\Delta}|Z_{j}|\leq c_{\Delta}(\beta)\right)=1-\beta.$ (7)
@@
$\mathcal{C}_{n}(x)=\left[\widehat{m}(x)-c_{\Delta}(\beta)(\tau_{\Delta}(x)n)^{-1/2},\widehat{m}(x)+c_{\Delta}(\beta)(\tau_{\Delta}(x)n)^{-1/2}\right]$
@@
$\liminf_{n\to\infty}\inf_{m\in\mathcal{H}^{\alpha}(C_{H})}\mathbb{P}\left(m(x)\in\mathcal{C}_{n}(x),\,\forall x\in\mathcal{X}\right)\geq 1-\beta.$
@@
$\max_{j=1,\ldots,\Delta}|Z_{j}|=\mathcal{O}_{\mathbb{P}}(\sqrt{2\log\Delta})=\mathcal{O}_{\mathbb{P}}(\sqrt{\log n})$ (9)
due to $\log\Delta\lesssim\log n$ which is implied by assumption (6). Therefore, $c_{\Delta}(\beta)\lesssim\sqrt{\log n}$. Moreover, the quantile $c_{\Delta}(\beta)$ can be explicitly represented in terms of the quantile function of the standard normal distribution since
$1-\beta=\mathbb{P}\left(\max_{j=1,\ldots,\Delta}|Z_{j}|\leq c_{\Delta}(\beta)\right)=\mathbb{P}\left(|Z_{1}|\leq c_{\Delta}(\beta)\right)^{\Delta}=(2\Phi(c_{\Delta}(\beta))-1)^{\Delta}.$
with cdf $\Phi$ of $\mathcal{N}(0,1)$. Consequently, $c_{\Delta}(\beta)=\Phi^{-1}(\frac{1}{2}(1+(1-\beta)^{1/\Delta}))$ is easy to compute.
@@
By (1) and (2) we obtain
$c_{X}\operatorname{vol}(A_{\Delta}(x))\leq p_{\Delta}(x)\leq C_{X}\operatorname{vol}(A_{\Delta}(x))$ (10)
@@
$\frac{c_{X}}{C_{\sigma^{2}}}\operatorname{vol}(A_{\Delta}(x))\leq\tau_{\Delta}(x)\leq\frac{C_{X}}{c_{\sigma^{2}}}\operatorname{vol}(A_{\Delta}(x)).$ (11)
Due to Assumption 2.1 we have $\operatorname{vol}(A_{\Delta}(x))=\mathcal{O}(\Delta^{-1})$ uniformly in $x\in\mathcal{X}$ and thus the diameter of the confidence band $\mathcal{C}_{n}$ is of order $\sqrt{(\log n)\Delta/n}$ uniformly on $\mathcal{X}$.
Example 2.3.
Let $\mathcal{X}=[0,1]^{p}$ and consider a partition based on an equidistant grid such that all cells $(A_{k})_{k=1,\dots,\Delta}$ are hypercubes of the same size. Denoting the edge length of the cubes by $\delta>0$, we have $\Delta=\delta^{-p}$, $\operatorname{diam}(A_{k})=\sqrt{p}\delta=\sqrt{p}\Delta^{-1/p}$ and $\operatorname{vol}(A_{k})=\delta^{p}=\Delta^{-1}$. In the homoscedastic case with constant $\sigma^{2}(\cdot)\equiv\sigma^{2}$ we obtain $\tau_{\Delta}(x)=p_{\Delta}(x)/\sigma^{2}$. For uniformly distributed covariates on $0,1]^{p}$ we have $p_{\Delta}(x)=\Delta^{-1}$ and in that case the confidence band simplifies to
$\mathcal{C}_{n}(x)=\Big[\widehat{m}(x)-c_{\Delta}(\beta)\sigma\sqrt{\frac{\Delta}{n}},\widehat{m}(x)+c_{\Delta}(\beta)\sigma\sqrt{\frac{\Delta}{n}}\Big].$
@@
$n/\Delta^{1+2\alpha/p}\to 0\text{ and }n/(\Delta n^{2/\nu})\to\infty,$
where $2/\nu$ is arbitrary small if all moments of $\varepsilon$ exist. Since the approximation error of the histogram estimator is of order $\Delta^{-\alpha/p}$ while the stochastic error term is of order $(\Delta/n)^{1/2}$, the first assumption is an undersmoothing condition. The second condition implies that the average number of observations in all cells goes to infinity such that the estimator is consistent. As $p$ increases, the assumptions become more restrictive which is the common curse of dimensionality problem for nonparametric function estimation. However, the assumptions are not contradictory, and for every dimension $p$ no bias reduction is needed.
The special case of confidence bands based on a histogram with a cubic partition is also covered by Cattaneo et al., (2020, Theorem 6.1 in combination with Remark 6.2). For general partitions in dimensions $p\geq 2$ Theorem 6.2 in Cattaneo et al., (2020) requires more regular regression functions compared to Theorem 2.2. More precisely, they need $\Delta^{3}/n\to 0$ (in our notation) instead of (6). An optimal choice $\Delta=n^{p/(2\alpha+p)}$ (up to logarithms, see Corollary 2.5) satisfies $\Delta^{3}/n\to 0$ only for $\alpha>p$. Cattaneo et al., (2020) also assume differentiability of the regression function. Theorem 2.2 does not require that the regression function’s and covariate density’s regularities have to grow with the dimension, which is a common restriction in the literature on nonparametric regression with multivariate covariates, see Proksch, (2016) and Chao et al., (2017) for confidence bands, and Dette and Derbort, (2001) and Neumeyer and Van Keilegom, (2010), among others, for different estimation and testing problems.
Typically, confidence bands regression functions rely on an extreme value distribution that is independent of $\Delta$. Based on Theorem 2.2 and the Gumbel extreme value distribution, we can derive an analogous result. Defining
$a_{\Delta}$ $=$ $\sqrt{2\log\Delta}\qquad\text{and}$
$b_{\Delta}$ $=$ $\sqrt{2\log\Delta}-\frac{\log(\log\Delta)+\log(4\pi)-\log(4)}{\sqrt{8\log\Delta}},$
the critical value $c_{\Delta}(\beta)$ can be approximated by
$\widetilde{c}_{\Delta}(\beta)=b_{\Delta}-\frac{\log(-\log(1-\beta))}{a_{\Delta}}.$ (12)
Let the confidence interval $\widetilde{\mathcal{C}}_{n}(x)$ be defined as $\mathcal{C}_{n}(x)$, but replacing $c_{\Delta}(\beta)$ by $\widetilde{c}_{\Delta}(\beta)$. The proof of the corollary is given in Section 6.
@@
Under the assumptions of Theorem 2.2 it holds that
$\liminf_{n\to\infty}\inf_{m\in\mathcal{H}^{\alpha}(C_{H})}\mathbb{P}\left(m(x)\in\widetilde{\mathcal{C}}_{n}(x),\,\forall x\in\mathcal{X}\right)=1-\beta.$
In contrast to the corollary Theorem 2.2 provides bands with explicit dependence on $\Delta$, and the accuracy of the bands does not depend on the rate of convergence to that limit distribution. Indeed, the approximation accuracy of $\widetilde{c}_{\Delta}(\beta)$ increases only slowly in the number of cells $\Delta$ and consistently overestimates $c_{\Delta}(\beta)$ as shown in Figure 1. Applying an approximation by the maximum of Gaussian random variables instead of a limit distribution is also suggested by Kreiß et al., (2026) for confidence bands for the spectral density in time series analysis.
[…]
Figure 1: Absolute error $\widetilde{c}_{\Delta}(\beta)-c_{\Delta}(\beta)$ and relative error $(\widetilde{c}_{\Delta}(\beta)-c_{\Delta}(\beta))/c_{\Delta}(\beta)$ for $\beta=0.05$ in dependence of $\Delta$.
From Theorem 2.2 we obtain a uniform rate of convergence of the histogram estimator. Then we obtain the following corollary whose proof is given in Section 6.
Corollary 2.5.
Let the partition satisfy Assumption 2.1. Assume $\sup_{x\in\mathcal{X}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$ and $\mathbb{E}\left[|\varepsilon|^{\nu}\right]<\infty$ for some $4\leq\nu\in\mathbb{N}$ such that $(\log n)^{5}\Delta/n^{1-2/\nu}\to 0$. Then we have
$\sup_{x\in\mathcal{X}}|\widehat{m}(x)-m(x)|=\mathcal{O}_{\mathbb{P}}\left(\Delta^{-\alpha/p}+\sqrt{(\log n)\Delta/n}\right).$
For $\nu>2+p/\alpha$ and $\Delta=\lfloor(n/\log n)^{p/(2\alpha+p)}\rfloor$ we obtain the uniform rate of convergence $\mathcal{O}_{\mathbb{P}}((n/\log n)^{-\alpha/(2\alpha+p)})$.
@@
We will now consider the case where the covariate distribution and the variance function of the noise are unknown. To this end, let $\widehat{\tau}_{\Delta}(x)$ be an estimator for $\tau_{\Delta}(x)$ from (8) and define the confidence band as
$\widehat{\mathcal{C}}_{n}(x)=\left[\widehat{m}(x)-c_{\Delta}(\beta)(\widehat{\tau}_{\Delta}(x)n)^{-1/2},\widehat{m}(x)+c_{\Delta}(\beta)(\widehat{\tau}_{\Delta}(x)n)^{-1/2}\right].$
@@
Under the assumptions of Theorem 2.2 and if
$\sup_{x\in\mathcal{X}}\left|\frac{\widehat{\tau}_{\Delta}(x)}{\tau_{\Delta}(x)}-1\right|=o_{\mathbb{P}}((\log n)^{-3/2}),$
@@
$\liminf_{n\to\infty}\inf_{m\in\mathcal{H}^{\alpha}(C_{H})}\mathbb{P}\left(m(x)\in\widehat{\mathcal{C}}_{n}(x),\,\forall x\in\mathcal{X}\right)\geq 1-\beta.$
The proof is again postponed to Section 6. To construct the estimator $\widehat{\tau}_{\Delta}(x)$, we first consider the homoscedastic case where $\tau_{\Delta}(x)=p_{\Delta}(x)/\sigma^{2}$ can be estimated by $\widehat{\tau}_{\Delta}(x)=\widehat{p}_{\Delta}(x)/\widehat{\sigma}^{2}$ with
$\widehat{p}_{\Delta}(x)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}.$ (13)
Note that $\widehat{p}_{\Delta}(x)$ is the density histogram estimator for the density function $f_{X}$ of the covariate distribution. This estimator has the desired rate which we show in Lemma 6.2. The constant variance can be estimated using residuals based on the histogram regression estimator, i.e.
$\widehat{\sigma}^{2}=\frac{1}{n}\sum_{i=1}^{n}\big(Y_{i}-\widehat{m}(X_{i})\big)^{2}.$